# The Catechol Benchmark: Time-series Solvent Selection Data for Few-shot Machine Learning

### v1

@@ 1 Introduction
Machine learning (ML) and artificial intelligence (AI) have showcased enormous potential in empowering the world of the natural sciences: from famous examples such as AlphaFold for protein predictions [1], to fusion reactor control [2], disease detection [3], battery design [4], and material discovery [5], among many more. However, we seldom see the machine learning community benchmark new methods in physical science datasets, mostly due to the difficulty in cleaning real-world data, the need for interdisciplinary understanding to correctly benchmark, and most importantly, how expensive the data can be to produce, resulting in many datasets being locked behind closed doors by large companies.
@@
Solvent selection is one of the biggest challenges for chemical manufacturing, with solvents often being the main source of waste in the manufacturing process [6]. Increased regulation on solvents and a drive to making process manufacturing more sustainable led to an interest in the discovery of greener solvents and for improved solvent replacement tools. However, most of the solvent replacement tools focus purely on learning unsupervised representations of solvents, with the hope that experimentalists can find solvents with similar properties to replace those with environmental concerns. A much stronger approach would consider the interaction of a variety of different solvents with a reaction of interest to directly predict reaction yields, in such a way that the best possible solvent can be selected according to a yield-sustainability trade-off.
Machine learning approaches have been shown to be a powerful tool for the prediction of chemical reaction conditions. Success has been reported in retro-synthesis [7, 8], condition recommendations [9], product predictions [10, 11], among others. While yield prediction has proven to be more difficult due to large inconsistencies in procedure and data reporting [12], we have still seen promising yield prediction results for smaller and more carefully curated datasets [13, 14, 15, 16]. However, these datasets lack the continuous reaction conditions, such as temperature and residence time, that are required to scale-up processes to practical manufacturing conditions.
In this paper, we release the first machine-learning-ready transient flow dataset, a framework that allows for quick and efficient screening of continuous reaction conditions. We specifically provide yield data over the uni-molecular allyl substituted catechol reaction, shown in Figure 1, with dense measurements across the residence time, temperature, and solvent space. We further showcase how this type of kinetic data poses new challenges to current machine learning methods for chemistry, and identify how the challenges can potentially be tackled by the community.
@@ 1.1 Related works
Reaction datasets are common in chemistry research, but their suitability for machine learning benchmarking tends to be poor. This can be a result of improper formatting or documentation, incomplete information about reaction conditions or the experimental set-up, or the lack of machine readability, leading to limited usage by the ML community. However, some effort has been made to address this, with the biggest example being the creation of the Open Reaction Database (ORD) [17], a repository containing over 2M different reactions, many of which come from US patent data (USPTO) [18]. However, the dataset falls short in some aspects, in particular with respect to machine learning readiness and data inconsistencies across reactions.
@@
ORDerly [12] allows for easy cleaning and preparation of ORD data, showing the promise of the dataset for forward and retro-synthetic prediction using transformers; however, it also shows that yield prediction cannot be done well due to data inconsistencies. Schwaller et al. [13] drew similar conclusions when using the USPTO dataset, stating that reaction conditions such as temperature, concentrations, and duration have a significant effect on yield. The assumption that every reaction in the dataset is optimized for reaction parameters proved too loose, resulting in inaccurate predictive models for yield, and highlighting the importance of creating datasets with full (including potentially sub-optimal) reaction conditions.
@@
More relevant to our work, Perera et al. [19] introduced a dataset of 5760 Suzuki-Miyaura cross-coupling reactions, Ahneman et al. [20] introduced a dataset of 3956 Buchwald–Hartwig aminations, and Prieto Kullmer et al. [21] investigated screening additives for Ni-catalysed reactions, all for the purposes of yield prediction. The datasets have been used in the benchmarking of Gaussian processes and Bayesian neural networks [14], deep learning models [13], language-model-based embeddings [16], data augmentation techniques [22], and Bayesian optimisation [15]. In each case, the datasets focus on discrete reaction variables, such as ligand, base, additives, or reactants at fixed temperatures and residence times. We are instead introducing a dataset rich in continuous reaction conditions (in our case temperature and residence time), as well as providing a pseudo-continuous representation of solvents themselves through the use of solvent mixtures.
@@
Perhaps the closest example to our dataset is presented in Nguyen et al. [23], who used high-throughput experimentation to screen 12708 catalyst informatics on the oxidative coupling of methane. In this case, they do provide process conditions for temperature, reactant equivalents, and flow rates; however, they do so in a discretized manner, as opposed to our approach that produces denser continuous representations of variables. The dataset has been used in the context of benchmarking language models for yield prediction, where the variables are used to create prompts to generate LLM embeddings of reactions. Introduced by Ramos et al. [24], the LLM embeddings are used for Bayesian optimization in reaction space. Recently, Ranković and Schwaller [16] fine tune a subset of 1180 LLM embeddings to use as deep kernel GPs, achieving even more favorable performance.
@@
More detailed and dense datasets including kinetic data usually have to be searched for in the research papers that originally published them, which are often accompanied by kinetic model fitting and benchmarking [25, 26, 27]. However, the datasets are seldom ML-ready, and tend to focus on variables which have predictable outcomes. In this work, we collect solvent data, which has a very large impact on the system’s dynamics and is often very challenging to model theoretically, making it a particularly interesting instance for machine learning applications.
@@ 1.2 The dataset
We further expand the dataset by including previous measurements on the same reaction class collected on a similar molecule, which was reported when developing the solvent ramping technology [28]. The two datasets are detailed in Table 1, and can be downloaded from Kaggle  [^1] Dataset: https://www.kaggle.com/datasets/aichemy/catechol-benchmark/.
@@
Allyl Substituted Catechol Solvent Mixtures 1227 24 SM + 2 Products
@@
Allyl Phenyl Ether Solvent Mixtures 283 11 SM + 1 Product $\times$
@@ 2.1 Transient flow and solvent ramping
Flow chemistry refers to a process in which the reaction is carried out in a continuous stream of reactant materials confined within tubing or narrow channels [29]. This technology offers an alternative to the traditional batch vessels typically used in chemical manufacturing and often presents benefits in areas such as safety, environmental impact, and scalability predictions [30]. Advantages include improved heat and mass transfer, more precise process control of the reaction conditions due to the smaller reaction volumes, and the ease of integration of online equipment and analytics [31].
Transient flow chemistry is an emerging technology used for collecting large quantities of reaction data in a continuous system. The method differs from traditional steady-state flow chemistry techniques, as the reaction conditions are varied continuously during experimentation to screen a range of reaction conditions [32]. Due to the efficient mixing inherent to flow systems, when the reaction conditions are adjusted in a controlled manner, each individual part of the flow is subject to different reaction conditions (i.e., plug flow), resulting in the efficient collection of a series of data [33]. An example of a relationship that can be investigated using this technique is the effect of the reaction time (i.e., the residence time of the reaction) on the yield of the reaction. This can be done by changing the cumulative flow rate in the reactor. The flow rates of the system are initially set to correspond to a specific residence time, and a step change to lower flow rates is performed. This means that the plugs at the end of the reactor experiences a shorter residence time in the reactor, and each subsequent volume of flow, or ‘plug,’ will have a longer residence time, producing a continuous data series of increasing residence times [34]. A visualization of this process is given in Figure 3.
@@
In a similar fashion, other variables can be investigated, such as temperature (varied by slowly ramping the reactor temperature) [35], and the equivalents of reagents in a reaction (by modifying the ratio of flow rates pumped from different reagent reservoirs) [36]. Reaction solvent is of particular interest in this dataset, following research interest in finding ‘green solvents’ as alternatives to traditional solvents [37, 28]. Solvent (mixtures) are treated as continuous variables, where the ratio of the two chosen solvents is varied using the ratio of the flow rates of the respective pumps to screen different solvent mixtures, and the changes in reaction yield are observed. Figure 2(a) shows all the solvent mixtures we were able to sample, as we effectively gather data between pure solvent pairs.
@@ 2.2 Solvent selection
In order to maximize the amount of information gathered from the data set, we used active learning. In particular, we trained a Gaussian process model on the Allyl Phenyl Ether dataset [28], which was the first published dataset that investigated solvent ramping using transient flow. We then selected a range of suitable available solvents [38] to create a set $\mathcal{S}$ and selected the solvents to query according to the entropy criterion:
@@ 2.3 Data acquisition and preprocessing
Online analytical measurements were collected using high-performance liquid chromatography (HPLC) [39], sampled every two minutes. This allows quantitative yield measurements to be collected over the course of the reaction, which can then be related to the reaction conditions applied to each sample. These can be calculated for each variable since we know the reactor volume, the flow rates, the temperature, and the duration that each sample in the flow stream was subjected to particular reaction conditions. The residence time is calculated by considering the measurement time, $t_{m}$, the reactor volume, $V$, and the cumulative flow rate function, $F_{c}(t)$. We then solve the equation:
[…]
@@ 3.1 Solvent featurization
As our goal is to predict yield surfaces on unseen solvents, a featurization that allows for transfer of information between solvents is required. Diorazio et al. [38] introduced a dataset of 272 solvents, with over 100 features for each, and further provide a 5-dimensional PCA representation of the solvent space. A second representation uses measurable properties of solvents [40], allowing easy grouping of solvents by type, e.g., as esters, ethers, and alkanes.
Cheminformatic features of molecules [41], ‘fragments’, are created using count vectors indicating the number of times a specific functional group appears in the molecule (following group contribution theory). Rogers and Hahn [42] show that bit vectors indicating the presence of substructures in a molecule, coined ‘molecular fingerprints’, can be used for molecular property prediction. We test the concatenation of both vectors, known as fragprints [43].
Finally, we investigate directly featurizing the reaction itself. This can be done, e.g., using the difference in sets containing molecular substructures in the starting materials and products [44, 45], known as differential reaction fingerprints (DRFP). Additionally, a reaction fingerprint can be learnt from larger open-source databases by using encoder-decoder neural networks [13], known as reaction fingerprints (RXNFP).
@@
A further question of interest is that of how to represent mixtures of solvents. Given a pair of solvents and their respective featurizations $S_{A}$ and $S_{B}$, we will initially take the naive approach of using a weighted mean: $S_{A\cup B}(b)=(1-b)S_{A}+bS_{B},$ where $b$ is the proportion of solvent B in the mixture. However, this linear transition can be an oversimplification of the underlying chemistry [46], so we investigate learning a non-linear mapping in Section 3.3.
@@ 3.2 Regression
When creating the test set, we take the mean of all repeated observations to avoid over-penalizing models that predict these reaction conditions poorly. We also omit the reactions containing acetic acid due to unintended a side-reaction [^2] The presence of acetic acid and high temperatures resulted in an unintended side-reaction of the expected product, possibly an esterification, as soon as it formed, showing very little yield of desired product but with high conversion. As such, we removed the affected results from our benchmark numbers.; creating models that are robust to unexpected side reactions poses an interesting future challenge.
Gaussian processes (GPs) are probabilistic, nonparametric models [47]. They are characterized by a covariance function, or kernel, that measures the similarity between a pair of inputs. These models provide uncertainty quantification, and perform well in low-data settings such as Bayesian optimization [48].
For our BaselineGP, we fit the data to the temperature and residence time inputs, ignoring the solvent, thus providing an improved proxy of one-hot encoding which has been shown to work well in low-data chemical regimes [49]. For the other GP methods, we transform the solvent into the featurized space, then use the Euclidean distance in that space as an input to an RBF kernel. In the appendix, we also evaluate a kernel over graphs, using pairwise shortest distances [50, 51].
Neural networks are a now-ubiquitous tool for regression. We start with a simple multi-layer perceptron (MLP). We also fine-tune pretrained large language models (LLMs): ChemBERTa [52], a model pretrained on chemistry texts, and RXNFP [13], pretrained on chemical reaction classifications. This follows works that have shown LLM abilities to generalize across reactions using their string representations [53]. In the appendix we further show results for neural process architectures which mimic GPs through meta-learning approaches [54].
Latent ODE methods use neural networks to model the latent state and dynamics of an ordinary differential equation (ODE) [55, 56], directly representing the underlying kinetics of the reaction. We also include an explicit state variant (LODE) [57], and a neural ODE (NODE) with time-dependent dynamics.
The regression results in Table 2 show the strength of the Spange featurization, which uses parameters known to be important to solvent effects. Whilst MLPs have a slight edge over GPs in their mean square error performance, the latter is also able to provide uncertainty quantification. Using LLM embeddings leads to poor performance, as reported previously, however, the same work reports strong performance when the embeddings are fine-tuned [16] which is left to future work.

### v2

@@ 1 Introduction
Machine learning (ML) and artificial intelligence (AI) have showcased enormous potential in empowering the world of the natural sciences: from famous examples such as AlphaFold for protein predictions [jumper2021highly], to fusion reactor control [degrave2022magnetic], disease detection [zhou2023foundation], battery design [folch2023combining], and material discovery [raccuglia2016machine], among many more. However, we seldom see the machine learning community benchmark new methods in physical science datasets, mostly due to the difficulty in cleaning real-world data, the need for interdisciplinary understanding to correctly benchmark, and most importantly, how expensive the data can be to produce, resulting in many datasets being locked behind closed doors by large companies.
@@
Solvent selection is one of the biggest challenges for chemical manufacturing, with solvents often being the main source of waste in the manufacturing process [constable2007perspective]. Increased regulation on solvents and a drive to making process manufacturing more sustainable led to an interest in the discovery of greener solvents and for improved solvent replacement tools. However, most of the solvent replacement tools focus purely on learning unsupervised representations of solvents, with the hope that experimentalists can find solvents with similar properties to replace those with environmental concerns. A much stronger approach would consider the interaction of a variety of different solvents with a reaction of interest to directly predict reaction yields, in such a way that the best possible solvent can be selected according to a yield-sustainability trade-off.
Machine learning approaches have been shown to be a powerful tool for the prediction of chemical reaction conditions. Success has been reported in retro-synthesis [karpov2019transformer, tetko2020state], condition recommendations [gao2018using], product predictions [coley2019graph, tu2022permutation], among others. While yield prediction has proven to be more difficult due to large inconsistencies in procedure and data reporting [wigh2024orderly], we have still seen promising yield prediction results for smaller and more carefully curated datasets [schwaller2021prediction, griffiths2023gauche, rankovic2024bayesian, rankovic2025gollum]. However, these datasets lack the continuous reaction conditions, such as temperature and residence time, that are required to scale-up processes to practical manufacturing conditions.
In this paper, we release the first machine-learning-ready transient flow dataset, a framework that allows for quick and efficient screening of continuous reaction conditions. We specifically provide yield data over the uni-molecular allyl substituted catechol reaction, shown in Figure 1, with dense measurements across the residence time, temperature, and solvent space. We answer the call for more flow chemistry reaction data [deadman2025wanted], further showcase how this type of kinetic data poses new challenges to current machine learning methods for chemistry, and identify potential solutions.
@@ 1.1 Related works
Reaction datasets are common in chemistry research, but their suitability for machine learning benchmarking tends to be poor. This can be a result of improper formatting or documentation, incomplete information about reaction conditions or the experimental set-up, or the lack of machine readability, leading to limited usage by the ML community. However, some effort has been made to address this, with the biggest example being the creation of the Open Reaction Database (ORD) [kearnes2021open], a repository containing over 2M different reactions, many of which come from US patent data (USPTO) [lowe2012extraction]. However, the dataset falls short in some aspects, in particular with respect to machine learning readiness and data inconsistencies across reactions.
@@
ORDerly [wigh2024orderly] allows for easy cleaning and preparation of ORD data, showing the promise of the dataset for forward and retro-synthetic prediction using transformers; however, it also shows that yield prediction cannot be done well due to data inconsistencies. schwaller2021prediction drew similar conclusions when using the USPTO dataset, stating that reaction conditions such as temperature, concentrations, and duration have a significant effect on yield. The assumption that every reaction in the dataset is optimized for reaction parameters proved too loose, resulting in inaccurate predictive models for yield, and highlighting the importance of creating datasets with full (including potentially sub-optimal) reaction conditions.
@@
More relevant to our work, perera2018platform introduced a dataset of 5760 Suzuki-Miyaura cross-coupling reactions, ahneman2018predicting introduced a dataset of 3956 Buchwald–Hartwig aminations, and prieto2022accelerating investigated screening additives for Ni-catalysed reactions, all for the purposes of yield prediction. The datasets have been used in the benchmarking of Gaussian processes and Bayesian neural networks [griffiths2023gauche], deep learning models [schwaller2021prediction], language-model-based embeddings [rankovic2025gollum], data augmentation techniques [schwaller2020data], and Bayesian optimisation [rankovic2024bayesian]. In each case, the datasets focus on discrete reaction variables, such as ligand, base, additives, or reactants at fixed temperatures and residence times. We are instead introducing a dataset rich in continuous reaction conditions (in our case temperature and residence time), as well as providing a pseudo-continuous representation of solvents themselves through the use of solvent mixtures.
@@
Perhaps the closest example to our dataset is presented in nguyen2019high, who used high-throughput experimentation to screen 12708 catalyst informatics on the oxidative coupling of methane. In this case, they do provide process conditions for temperature, reactant equivalents, and flow rates; however, they do so in a discretized manner, as opposed to our approach that produces denser continuous representations of variables. The dataset has been used in the context of benchmarking language models for yield prediction, where the variables are used to create prompts to generate LLM embeddings of reactions. Introduced by ramos2023bayesian, the LLM embeddings are used for Bayesian optimization in reaction space. Recently, rankovic2025gollum fine tune a subset of 1180 LLM embeddings to use as deep kernel GPs, achieving even more favorable performance.
@@
More detailed and dense datasets including kinetic data usually have to be searched for in the research papers that originally published them, which are often accompanied by kinetic model fitting and benchmarking [sagmeister2023accelerating, sagmeister2024simultaneous, silber2023accelerating]. However, the datasets are seldom ML-ready, and tend to focus on variables which have predictable outcomes. In this work, we collect solvent data, which has a very large impact on the system’s dynamics and is often very challenging to model theoretically, making it a particularly interesting instance for machine learning applications.
@@ 1.2 The dataset
We further expand the dataset by including previous measurements on the same reaction class collected on a similar molecule, which was reported when developing the solvent ramping technology [WO2025073762]. The two datasets are detailed in Table 1, and can be downloaded from Kaggle  [^1] Dataset: https://www.kaggle.com/datasets/aichemy/catechol-benchmark/.
@@
. We further include sustainability details for all the solvents screened in Table 2 according to the GSK guide [henderson2011expanding].
@@
Allyl Substituted
Catechol
Solvent Mixtures 1227 24 SM + 2 Products
@@
Allyl Phenyl
Ether
Solvent Mixtures 283 11 SM + 1 Product $\times$
Table 2: List of screened solvents and their classification. For more details see henderson2011expanding.
Classification Solvents
Green Ethylene Glycol; IPA; Water; Ethanol; Cyrene; Ethyl Acetate; DMC
Situation dependent Methanol; 2-MeTHF; Cyclohexane; Acetonitrile; Acetic Acid; 2-Butanone; t-Butanol
Needs replacement Diethyl Ether; DMA; THF; MTBE
Not on GSK guide HFIP; 2,2,2-TFE; Decanol; Methyl Propionate; Ethyl $\ell$-Lactate
@@ 2.1 Transient flow and solvent ramping
Flow chemistry refers to a process in which the reaction is carried out in a continuous stream of reactant materials confined within tubing or narrow channels [plutschack2017hitchhiker]. This technology offers an alternative to the traditional batch vessels typically used in chemical manufacturing and often presents benefits in areas such as safety, environmental impact, and scalability predictions [hartman2011deciding]. Advantages include improved heat and mass transfer, more precise process control of the reaction conditions due to the smaller reaction volumes, and the ease of integration of online equipment and analytics [boyall2024automated].
Transient flow chemistry is an emerging technology used for collecting large quantities of reaction data in a continuous system. The method differs from traditional steady-state flow chemistry techniques, as the reaction conditions are varied continuously during experimentation to screen a range of reaction conditions [williams2024dynamic]. Due to the efficient mixing inherent to flow systems, when the reaction conditions are adjusted in a controlled manner, each individual part of the flow is subject to different reaction conditions (i.e., plug flow), resulting in the efficient collection of a series of data [schrecker2023discovery]. An example of a relationship that can be investigated using this technique is the effect of the reaction time (i.e., the residence time of the reaction) on the yield of the reaction. This can be done by changing the cumulative flow rate in the reactor. The flow rates of the system are initially set to correspond to a specific residence time, and a step change to lower flow rates is performed. This means that the plugs at the end of the reactor experiences a shorter residence time in the reactor, and each subsequent volume of flow, or ‘plug,’ will have a longer residence time, producing a continuous data series of increasing residence times [schrecker2024comparative]. A visualization of this process is given in Figure 3.
@@
In a similar fashion, other variables can be investigated, such as temperature (varied by slowly ramping the reactor temperature) [moore2016kinetics], and the equivalents of reagents in a reaction (by modifying the ratio of flow rates pumped from different reagent reservoirs) [mcmullen2023automated]. Reaction solvent is of particular interest in this dataset, following research interest in finding ‘green solvents’ as alternatives to traditional solvents [drelinkiewicz2024rapid, WO2025073762]. Solvent (mixtures) are treated as continuous variables, where the ratio of the two chosen solvents is varied using the ratio of the flow rates of the respective pumps to screen different solvent mixtures, and the changes in reaction yield are observed. Figure 2(a) shows all the solvent mixtures we were able to sample, as we effectively gather data between pure solvent pairs.
@@ 2.2 Solvent selection
In order to maximize the amount of information gathered from the data set, we used active learning. In particular, we trained a Gaussian process model on the Allyl Phenyl Ether dataset [WO2025073762], which was the first published dataset that investigated solvent ramping using transient flow. We then selected a range of suitable available solvents [diorazio2016toward] to create a set $\mathcal{S}$ and selected the solvents to query according to the entropy criterion:
@@ 2.3 Data acquisition and preprocessing
Online analytical measurements were collected using high-performance liquid chromatography (HPLC) [dixon2024operator], sampled every two minutes. This allows quantitative yield measurements to be collected over the course of the reaction, which can then be related to the reaction conditions applied to each sample. These can be calculated for each variable since we know the reactor volume, the flow rates, the temperature, and the duration that each sample in the flow stream was subjected to particular reaction conditions. The residence time is calculated by considering the measurement time, $t_{m}$, the reactor volume, $V$, and the cumulative flow rate function, $F_{c}(t)$. We then solve the equation:
$V=\int_{t_{i}}^{t_{m}}F_{c}(t)dt$ (2)
@@ 3.1 Solvent featurization
As our goal is to predict yield surfaces on unseen solvents, a featurization that allows for transfer of information between solvents is required. diorazio2016toward introduced a dataset of 272 solvents, with over 100 features for each, and further provide a 5-dimensional PCA representation of the solvent space. A second representation uses measurable properties of solvents [spange2021reappraisal], allowing easy grouping of solvents by type, e.g., as esters, ethers, and alkanes.
Cheminformatic features of molecules [landrum2006rdkit], ‘fragments’, are created using count vectors indicating the number of times a specific functional group appears in the molecule (following group contribution theory). rogers2010extended show that bit vectors indicating the presence of substructures in a molecule, coined ‘molecular fingerprints’, can be used for molecular property prediction. We test the concatenation of both vectors, known as ‘fragprints’ [griffiths2022data].
Finally, we investigate directly featurizing the reaction itself. This can be done, e.g., using the difference in sets containing molecular substructures in the starting materials and products [schneider2015development, probst2022reaction], known as differential reaction fingerprints (DRFP). Additionally, a reaction fingerprint can be learnt from larger open-source databases by using encoder-decoder neural networks [schwaller2021prediction], known as reaction fingerprints (RXNFP).
@@
A further question of interest is that of how to represent mixtures of solvents. Given a pair of solvents and their respective featurizations $S_{A}$ and $S_{B}$, we will initially take the naive approach of using a weighted mean: $S_{A\cup B}(b)=(1-b)S_{A}+bS_{B},$ where $b$ is the proportion of solvent B in the mixture. However, this linear transition can be an oversimplification of the underlying chemistry [burrows2022solvation], so we investigate learning a non-linear mapping in Section 3.3.
@@ 3.2 Regression
When creating the test set, we take the mean of all repeated observations to avoid over-penalizing models that predict these reaction conditions poorly. We also omit the reactions containing acetic acid due to unintended a side-reaction [^2] The presence of acetic acid and high temperatures resulted in an unintended side-reaction of the expected product - possibly an esterification - as soon as it formed, showing very little yield of desired product but with high conversion. As such, we removed the affected results from our benchmark numbers.; creating models that are robust to unexpected side reactions poses an interesting future challenge.
Gaussian processes (GPs) are probabilistic, nonparametric models [rasmussen2005gaussian]. They are characterized by a covariance function, or kernel, that measures the similarity between a pair of inputs. These models provide uncertainty quantification, and perform well in low-data settings such as Bayesian optimization [frazier2018tutorial].
For our BaselineGP, we fit the data to the temperature and residence time inputs, ignoring the solvent, thus providing an improved proxy of one-hot encoding which has been shown to work well in low-data chemical regimes [pomberger2022effect]. For the other GP methods, we transform the solvent into the featurized space, then use the Euclidean distance in that space as an input to an RBF kernel. In the appendix, we also evaluate a kernel over graphs, using pairwise shortest distances [borgwardt2005shortest, xie2025bogrape, xie2025global].
Neural networks are a now-ubiquitous tool for regression. We start with a simple multi-layer perceptron (MLP). We also fine-tune pretrained large language models (LLMs): ChemBERTa [chithrananda2020chemberta], a model pretrained on chemistry texts, and RXNFP [schwaller2021prediction], pretrained on chemical reaction classifications. This follows works that have shown LLM abilities to generalize across reactions using their string representations [bran2024transformers]. In the appendix we further show results for neural process architectures which mimic GPs through meta-learning approaches [garnelo2018conditional].
Latent ODE methods use neural networks to model the latent state and dynamics of an ordinary differential equation (ODE) [chen2018neural, qing2024system], directly representing the underlying kinetics of the reaction. We also include an explicit state variant (LODE) [rubanova2019latent], and a neural ODE (NODE) with time-dependent dynamics.
The regression results in Table 3 show the strength of the Spange featurization, which uses parameters known to be important to solvent effects. Whilst MLPs have a slight edge over GPs in their mean square error performance, the latter is also able to provide uncertainty quantification. Using LLM embeddings leads to poor performance, as reported previously, however, the same work reports strong performance when the embeddings are fine-tuned [rankovic2025gollum] which is left to future work.